The dimensions of a classroom are 10 m, 8 m and 5 m.the cost of decorating it's walls by a wallpaper of 1 m width and which costs rs.50 per meter is -----?
Answer by ewatrrr(24785)

 
Hi,
.
Surface Area(walls) = 2(10m•5m) + 2(8m•5m) = 180m^2
1m•180m = 180m^2
rs.50*180 = rs90